OpenStudy (technodynamic):

Let v(t)=t^2−3t be the velocity, in feet per second, of an object at time t, in seconds. a) What is the initial velocity? ___ ft/sec b) When does the object have a velocity of zero? If there is more than time, list all answers in a comma separated list. ___ seconds

OpenStudy (technodynamic):

It's calculus, but You're right I should try there too. thank you very much.

OpenStudy (anonymous):

initial velocity means v(0) so 0.. velocity of zero so equate v(t) to zero 0 = t^2 - 3t t(t-3) = 0 at t= 0 (as we found at a) and t = 3